BIO 340 ACTUAL STUDY GUIDE ANSWERS AND
QUESTIONS SET A+
✔✔In the biosynthetic pathway for conversion from homoserine to methionine, you
identify a double mutant Met1/Met2. This mutant will grow only if which supplement(s)
are added to minimal media? - ✔✔Methionine only
✔✔You have 2 true-breeding strains of lucky cat in your lab. The W strain has white fur
and the G strain has golden fur. You cross W and G strains. The F1 from this cross
have golden fur. When you cross 2 F1's, the F2's have either golden or white fur. In the
F2's, the ratio of golden to white is 9:1. Develop a hypothesis to explain this ratio. -
✔✔Recessive lethal
The allele can be both dominant (that's why the F1 generation all has golden fur) AND
recessive lethal (that's why the F2 generation has both white and golden fur in a 9:1
ratio)
✔✔The O locus in domesticated cats is X-linked. The big O allele produces orange
coloration and the little o allele produces black coloration. Males are hemizygous for the
locus and females are mosaics; therefore, the dominance relationship between the
alleles cannot be determined by crossing experiments. Describe a possible research
strategy to determine the dominance relationship between the big O and little o alleles. -
✔✔
✔✔A new species of grass has been identified with 2 genes located 10 cM apart. Gene
A produces a hormone that causes stems to grow. Gene B produces small seeds. Both
genes show complete dominance over their recessive forms. A pure-breeding plant with
a long stem and large seeds is crossed with a pure-breeding plant with a short stem and
small seeds. The F1 offspring are then crossed with each other or produce the F2
generation. What are the expected phenotypic ratios of the F2 generation? - ✔✔Parent
Genotypes: AAbb x aaBB
(Recombinants are AB and ab)
10 cM = 0.1 crossover rate (10%)
,Because AB and ab are the recombinants, they each have a 5% chance of being
passed on. This makes the probability that the gametes Ab and aB have a 45% chance
each of being passed on.
Ab = 45%
aB = 45%
AB = 5%
ab = 5%
Possible Combos of F2 generation:
Ab/aB = .45 x .45 = 0.2025
Ab/AB = .45 x .05 = 0.0225
Ab/ab = .45 x .05 = 0.0225
Ab/Ab = .45 x .45 = 0.2025
aB/aB = .45 x .45 = 0.2025
aB/Ab = .45 x .45 = 0.2025
aB/AB = .45 x .05 = 0.0225
aB/ab = .45 x .05 = 0.0225
AB/AB = .05 x .05 = 0.0025
AB/Ab = .05 x .45 = 0.0225
AB/aB = .05 x .45 = 0.0225
AB/ab = .05 x .05 = 0.0025
ab/ab = .05 x .05 = 0.0025
ab/Ab = .05 x .45 = 0.0225
ab/aB = .05 x .45 = 0.0225
ab/AB = .05 x .05 = 0.0025
Grown Stem, Small Seed (A-B-) = 0.5025
Non-grown Stem, Big Seed (aabb) = 0.0025
Grown Stem, Big Seed (A-bb) = 0.2475
Non-grown Stem, Small Seed (aaB-) = 0.2475
✔✔Titin (T) is the largest gene in the cattle genome. Your herd contains 3 deleterious
mutations: T1 is dominant and located at 1 cM from the start of the gene. T2 is
recessive and located at 3 cM from the start. T3 is also recessive and located 5 cM from
the start. Your bull carries both recessive and mutations on separate chromosomes,
while your cow is heterozygous for the dominant mutation. What is the probability that
when mated they produce an offspring that is not a carrier for any of the mutations. -
✔✔(0.5)(0.5)(0.02) = 0.005
0.5 chance of getting fully functional allele (+++) from mom, 0.5 chance of getting ++-
from dad, and 2% chance crossover.
0.05 - 0.03 = 0.02 (how far apart they are from each other)
,Bull: ++(T3)/+(T2)+
Cow: +++/(T1)++
✔✔A troop of Japanese snow monkeys have become a nuisance in their native, wintery
range. As part of an experiment, they have been relocated to Apache Junction, AZ.
Within the troop, there are phenotypic differences for both fur length and fur density, but
there are no genetic differences for these traits. Would you expect the monkey troop to
evolve in response to their new environment, with respect of these phenotypes? Why or
why not? - ✔✔No, because there is no genetic basis. Therefore, traits would not be
passed down.
✔✔Parent-child and full-sibling kinship coefficients are both 25%. By looking at the
patterns of IBD in their genomes, how can you distinguish the relationships? -
✔✔Parent-child will have an entire chromosome in common, while only full siblings
have parts of chromosomes in common.
✔✔Many plants exhibit a mating system known as mixed-mating. A certain fraction (s)
of seeds are produced by selfing and the rest (1-s) by outcrossing. Assuming that a
plant population is not currently inbred, after one generation, what fraction of the
population is autozygous? - ✔✔0.5s (half of the selfing seeds will be autozygous)
1/N seeds are a result of selfing and 1/2N will be autozygous
✔✔Monkeys infected with Devil Pox survive depending on their genotype at a single
locus. A1A1 individuals survive 50% of the time, A1A2 individuals survive 75% of the
time, and A2A2 individuals survive 25% of the time. A troop of 10 A1A1, 12 A1A2, and 4
A2A2 juveniles are infected. What are the allele frequencies after one generation of
viability selection? - ✔✔A1A1: 10 x 0.5 = 5
A1A2: 12 x 0.75 = 9
A2A2: 4 x 0.25 = 1
5 + 9 + 1 = 15
A1 Frequency = [(5 x 2) + 9]/(2 x 15) = 19/(2 x 15) = 0.63
A2 Frequency = [(1 x 2) + 9]/(2 x 15) = 11/(2 x 15) = 0.36
Have to multiply by 2 because there are 2 copies of that particular allele, etc.
✔✔Wingless (wg) is typically a deleterious mutation in Drosophila. However, under
some environmental conditions it can be favored. On the windy shores of the British
Coast, the fitness of genotypes are as follows:
*wg/wg = 1.2
*+/wg = 1.1
*+/+ = 1
, What is the probability of fixation of a novel wg mutation that arises in a colony of 50
British coastal flies with stable population size? - ✔✔The probability of fixation is 1/2N
1/2N = 1/2(50) = 0.01 = 1%
✔✔Probability - ✔✔the frequency of that event among all possible events
exp. tossing a coin
✔✔Joint Probability - ✔✔the probability that two events occur together
✔✔Additive Rule of Probability - ✔✔if two events, A and B, are mutually exclusive, then
the probability that either one occurs is the sum of their individual probabilities
✔✔Multiplicative Rule of Probability - ✔✔If two events are independent, then their joint
probability is given by:
P(A and B)= P(A) x P(B)
✔✔Your bike has a flat tire with the probability 1/10. Your car has a flat tire with a
probability 1/100. Your care and bike both have flat tires with probability 1/110. Are
these events independent? - ✔✔False
✔✔Mendel's five critical Experimental Innovations - ✔✔1. controlled crosses between
plants
2. use of true-breeding strains
3. selection of dichotomous traits
4. quantification of results
5. use of replicate, reciprocal, and test crosses
✔✔True-Breeding Strains - ✔✔a group of identical individuals that always produce
offspring of the same phenotype when mated with each other
✔✔The true breeding P1 are - ✔✔homozygous
✔✔The purple F2 are - ✔✔heterozygous
✔✔The purple and white F2 are - ✔✔heterozygous and homozygous
✔✔Mendel's Law of Segregation - ✔✔a. the two copies of a gene separate equally
during gamete formation, so that half of the gametes carry one member of the pair, and
half carry the other
b. during fertilization, gametes unite at random, i.e. they fuse without regard to which
allele is carried by the other one
QUESTIONS SET A+
✔✔In the biosynthetic pathway for conversion from homoserine to methionine, you
identify a double mutant Met1/Met2. This mutant will grow only if which supplement(s)
are added to minimal media? - ✔✔Methionine only
✔✔You have 2 true-breeding strains of lucky cat in your lab. The W strain has white fur
and the G strain has golden fur. You cross W and G strains. The F1 from this cross
have golden fur. When you cross 2 F1's, the F2's have either golden or white fur. In the
F2's, the ratio of golden to white is 9:1. Develop a hypothesis to explain this ratio. -
✔✔Recessive lethal
The allele can be both dominant (that's why the F1 generation all has golden fur) AND
recessive lethal (that's why the F2 generation has both white and golden fur in a 9:1
ratio)
✔✔The O locus in domesticated cats is X-linked. The big O allele produces orange
coloration and the little o allele produces black coloration. Males are hemizygous for the
locus and females are mosaics; therefore, the dominance relationship between the
alleles cannot be determined by crossing experiments. Describe a possible research
strategy to determine the dominance relationship between the big O and little o alleles. -
✔✔
✔✔A new species of grass has been identified with 2 genes located 10 cM apart. Gene
A produces a hormone that causes stems to grow. Gene B produces small seeds. Both
genes show complete dominance over their recessive forms. A pure-breeding plant with
a long stem and large seeds is crossed with a pure-breeding plant with a short stem and
small seeds. The F1 offspring are then crossed with each other or produce the F2
generation. What are the expected phenotypic ratios of the F2 generation? - ✔✔Parent
Genotypes: AAbb x aaBB
(Recombinants are AB and ab)
10 cM = 0.1 crossover rate (10%)
,Because AB and ab are the recombinants, they each have a 5% chance of being
passed on. This makes the probability that the gametes Ab and aB have a 45% chance
each of being passed on.
Ab = 45%
aB = 45%
AB = 5%
ab = 5%
Possible Combos of F2 generation:
Ab/aB = .45 x .45 = 0.2025
Ab/AB = .45 x .05 = 0.0225
Ab/ab = .45 x .05 = 0.0225
Ab/Ab = .45 x .45 = 0.2025
aB/aB = .45 x .45 = 0.2025
aB/Ab = .45 x .45 = 0.2025
aB/AB = .45 x .05 = 0.0225
aB/ab = .45 x .05 = 0.0225
AB/AB = .05 x .05 = 0.0025
AB/Ab = .05 x .45 = 0.0225
AB/aB = .05 x .45 = 0.0225
AB/ab = .05 x .05 = 0.0025
ab/ab = .05 x .05 = 0.0025
ab/Ab = .05 x .45 = 0.0225
ab/aB = .05 x .45 = 0.0225
ab/AB = .05 x .05 = 0.0025
Grown Stem, Small Seed (A-B-) = 0.5025
Non-grown Stem, Big Seed (aabb) = 0.0025
Grown Stem, Big Seed (A-bb) = 0.2475
Non-grown Stem, Small Seed (aaB-) = 0.2475
✔✔Titin (T) is the largest gene in the cattle genome. Your herd contains 3 deleterious
mutations: T1 is dominant and located at 1 cM from the start of the gene. T2 is
recessive and located at 3 cM from the start. T3 is also recessive and located 5 cM from
the start. Your bull carries both recessive and mutations on separate chromosomes,
while your cow is heterozygous for the dominant mutation. What is the probability that
when mated they produce an offspring that is not a carrier for any of the mutations. -
✔✔(0.5)(0.5)(0.02) = 0.005
0.5 chance of getting fully functional allele (+++) from mom, 0.5 chance of getting ++-
from dad, and 2% chance crossover.
0.05 - 0.03 = 0.02 (how far apart they are from each other)
,Bull: ++(T3)/+(T2)+
Cow: +++/(T1)++
✔✔A troop of Japanese snow monkeys have become a nuisance in their native, wintery
range. As part of an experiment, they have been relocated to Apache Junction, AZ.
Within the troop, there are phenotypic differences for both fur length and fur density, but
there are no genetic differences for these traits. Would you expect the monkey troop to
evolve in response to their new environment, with respect of these phenotypes? Why or
why not? - ✔✔No, because there is no genetic basis. Therefore, traits would not be
passed down.
✔✔Parent-child and full-sibling kinship coefficients are both 25%. By looking at the
patterns of IBD in their genomes, how can you distinguish the relationships? -
✔✔Parent-child will have an entire chromosome in common, while only full siblings
have parts of chromosomes in common.
✔✔Many plants exhibit a mating system known as mixed-mating. A certain fraction (s)
of seeds are produced by selfing and the rest (1-s) by outcrossing. Assuming that a
plant population is not currently inbred, after one generation, what fraction of the
population is autozygous? - ✔✔0.5s (half of the selfing seeds will be autozygous)
1/N seeds are a result of selfing and 1/2N will be autozygous
✔✔Monkeys infected with Devil Pox survive depending on their genotype at a single
locus. A1A1 individuals survive 50% of the time, A1A2 individuals survive 75% of the
time, and A2A2 individuals survive 25% of the time. A troop of 10 A1A1, 12 A1A2, and 4
A2A2 juveniles are infected. What are the allele frequencies after one generation of
viability selection? - ✔✔A1A1: 10 x 0.5 = 5
A1A2: 12 x 0.75 = 9
A2A2: 4 x 0.25 = 1
5 + 9 + 1 = 15
A1 Frequency = [(5 x 2) + 9]/(2 x 15) = 19/(2 x 15) = 0.63
A2 Frequency = [(1 x 2) + 9]/(2 x 15) = 11/(2 x 15) = 0.36
Have to multiply by 2 because there are 2 copies of that particular allele, etc.
✔✔Wingless (wg) is typically a deleterious mutation in Drosophila. However, under
some environmental conditions it can be favored. On the windy shores of the British
Coast, the fitness of genotypes are as follows:
*wg/wg = 1.2
*+/wg = 1.1
*+/+ = 1
, What is the probability of fixation of a novel wg mutation that arises in a colony of 50
British coastal flies with stable population size? - ✔✔The probability of fixation is 1/2N
1/2N = 1/2(50) = 0.01 = 1%
✔✔Probability - ✔✔the frequency of that event among all possible events
exp. tossing a coin
✔✔Joint Probability - ✔✔the probability that two events occur together
✔✔Additive Rule of Probability - ✔✔if two events, A and B, are mutually exclusive, then
the probability that either one occurs is the sum of their individual probabilities
✔✔Multiplicative Rule of Probability - ✔✔If two events are independent, then their joint
probability is given by:
P(A and B)= P(A) x P(B)
✔✔Your bike has a flat tire with the probability 1/10. Your car has a flat tire with a
probability 1/100. Your care and bike both have flat tires with probability 1/110. Are
these events independent? - ✔✔False
✔✔Mendel's five critical Experimental Innovations - ✔✔1. controlled crosses between
plants
2. use of true-breeding strains
3. selection of dichotomous traits
4. quantification of results
5. use of replicate, reciprocal, and test crosses
✔✔True-Breeding Strains - ✔✔a group of identical individuals that always produce
offspring of the same phenotype when mated with each other
✔✔The true breeding P1 are - ✔✔homozygous
✔✔The purple F2 are - ✔✔heterozygous
✔✔The purple and white F2 are - ✔✔heterozygous and homozygous
✔✔Mendel's Law of Segregation - ✔✔a. the two copies of a gene separate equally
during gamete formation, so that half of the gametes carry one member of the pair, and
half carry the other
b. during fertilization, gametes unite at random, i.e. they fuse without regard to which
allele is carried by the other one